OpenStudy (anonymous):

Consider the line y=-9x+9 Find the equations of the parallel & perpendicular lines that pass through (2,6)

OpenStudy (anonymous):

i think parallel is y=-9x+24 & perpendicular is y=1/9x+52/9 is this correct?

OpenStudy (anonymous):

The only thing that changes is the y-intercept y=mx+b Parallel: Keep the the slope the same y=-9x+b The point 2,6 is a solution to that equation so we can plug that in too. 6=-9(2)+b Now, solve for b. Perpendicular: slope is opposite reciprocal so, 1/6 is our slope y=1/6x+b Then plug in the stuff that you know which is the point 2,6 6=(1/6)2+b Solve for b.

OpenStudy (anonymous):

can you read my first post? i have the answers.. i just want to know if they are correct!

OpenStudy (anonymous):

Parallel is correct. Perpendicular is not.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

Multiplying (1/6)2 is 1/3 Did you get that?

OpenStudy (anonymous):

it wouldnt be 1/6.. it would be 1/9..

OpenStudy (anonymous):

sorry, i mean 2*1/6 which is 2/6 simplified to 1/3